# Mixed-Precision Quantization: Make the Best Use of Bits Where They Matter Most

### v3

@@ II-D Convergence Analysis
Given $10<w<c+$ 1 and eigenvalue $\lambda_{\max}({\bf P})<\frac{1}{2\sqrt{c^{2}+w^{2}}}$, the dynamic system described by (10) and (11) converges to an equilibrium point, where
@@
Theorem 1 reveals that the PPSO and GC-PSO algorithm will converge to an equilibrium point regardless of the exact value of $c$ and $w$, provided that $10<w<c+$ 1 and $\lambda_{\max}({\bf P})<\frac{1}{2\sqrt{c^{2}+w^{2}}}$. It should also be pointed out that the convergence condition is derived based on the worst-case analysis, which may be conservative.
@@ III-A Problem Statement
Figure 3: The illustration of direct FIR filter.
As shown in Fig. 3, we consider an $N$-tap linear-phase direct FIR filter with real-valued impulse response ${\bf h}=\{h[n]\}_{n=0}^{N-1}$. The corresponding frequency response is given by
@@ V-C1 Least Squares
The convergence performance of the proposed PPSO and GC-PSO algorithms is measured by the error term $\left\|\mathbf{z}_{t}-\mathbf{z}^{*}\right\|_{2}$, where $\mathbf{z}_{t}$ is the computed parameter at the end of $t$-th iteration of quantized GD. As shown in Fig. 9a and 9b, the proposed PPSO and GC-PSO algorithms achieve faster convergence than naive rounding (fixed-precision quantization).

### v4

@@ II-D Convergence Analysis
Given $10<w<c+$ 1 and the largest eigenvalue of matrix $\mathbf{P}$ $\lambda_{\max}({\bf P})<\frac{1}{2\sqrt{c^{2}+w^{2}}}$, the dynamic system described by (10) and (11) converges to an equilibrium point, where
@@
Theorem 1 presents a sufficient condition for the convergence of the proposed PPSO and GC-PSO algorithms. It is derived by modeling the simplified deterministic dynamics of particle motion as a discrete-time system. The result shows that under $10<w<c+$ 1 and $\lambda_{\max}({\bf P})<\frac{1}{2\sqrt{c^{2}+w^{2}}}$, the particles will converge to a stable equilibrium point, despite the non-linear rounding operation. It should also be pointed out that the convergence condition is derived based on the worst-case analysis, which may be conservative.
@@ III-A Problem Statement
We consider an $N$-tap linear-phase direct FIR filter with real-valued impulse response ${\bf h}=\{h[n]\}_{n=0}^{N-1}$. The corresponding frequency response is given by
@@ V-C1 Least Squares
The convergence performance of the proposed PPSO and GC-PSO algorithms is measured by the error term $\left\|\mathbf{z}_{t}-\mathbf{z}^{*}\right\|_{2}$, where $\mathbf{z}_{t}$ is the computed parameter at the end of $t$-th iteration of quantized GD. As shown in Fig. 8a and 8b, the proposed PPSO and GC-PSO algorithms achieve faster convergence than naive rounding (fixed-precision quantization).